OpenStudy (emilyjones284):

MEDAL AND FAN! (: Find the probability distribution for the following function:

OpenStudy (emilyjones284):

\[P(x) =\frac{ x }{ 2 } \] for x = 4, 6, and 8

jimthompson5910 (jim_thompson5910):

when x = 4, what is P(x)

jimthompson5910 (jim_thompson5910):

hmm this function is a bit off

jimthompson5910 (jim_thompson5910):

are there any typos?

OpenStudy (emilyjones284):

ohhhh im sorry its \[\frac{ x }{ 2 } +1\]

jimthompson5910 (jim_thompson5910):

that's still looks off what's supposed to happen is that P(x) should be a small number between 0 and 1 for each value of x

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the problem?

jimthompson5910 (jim_thompson5910):

I'm thinking it's more like 2/x, but I don't know really it could be anything

OpenStudy (emilyjones284):

yeah hold on i will

jimthompson5910 (jim_thompson5910):

ok thanks

OpenStudy (emilyjones284):

sorry its taking so long itll just be a second

jimthompson5910 (jim_thompson5910):

its ok

jimthompson5910 (jim_thompson5910):

hmm that's annoying and a bad use of notation

jimthompson5910 (jim_thompson5910):

oh well

jimthompson5910 (jim_thompson5910):

anyways, P(x) = (x/2) + 1 P(4) = (4/2) + 1 P(4) = 2 + 1 P(4) = 3

jimthompson5910 (jim_thompson5910):

the same is done for the other x values

OpenStudy (emilyjones284):

ohhhh okay thats actually pretty easy haha thanks (:

jimthompson5910 (jim_thompson5910):

yeah it's not too bad, but they're using really odd notation for probabilities (since probabilities should be between 0 and 1)